OpenStudy (anonymous):

Playing shortstop, you pick up a ground ball and throw it to second base. The ball is thrown horizontally, with a speed of 25m/s , directly toward point A . When the ball reaches the second baseman 0.41s later, it is caught at point B. What is the distance of vertical drop, AB?

OpenStudy (anonymous):

lets start with drawing to make obvious  Calculate V y component by this formula \[x = 0.5 a t ^{2} = 0.5* 9.8 * 0.41^{2}= 0.82369 m.\] Calculate V x component ( it has nothing to affect it as we neglected air resistance). \[y = Vt = 25* 041= 10.25 m.\] and now by Pythagoras theorem. \[ab = \sqrt{x ^{2}+y ^{2}}=\sqrt{10.25^{2}+0.82369^{2}}= 10.2830426 \approx 10.3 m\] There you go.I hope you didn't get confused from distance x and y as I drew the reversely , so I used them as I drew.